# Malle’s conjecture and Brauer groups of stacks

### v2

@@ 3.2. Orbifold Picard group of $BG$
Condition (2) says that $\zeta^{w(\gamma)}=\chi(\gamma(\zeta))$ for all $\gamma\in G(-1)(k^{\mathrm{sep}})$ and all $\zeta\in\mu_{n}$ with $n$ not divisible by the characteristic of $k$; via a choice of isomorphism from Lemma 3.3 this becomes $\zeta^{w(g)}=\chi(g)$ for all $g\in G(k^{\mathrm{sep}})$ and all such $n$. In particular if $\chi$ is trivial then $w$ takes integer values.
@@
If $w$ is $\widehat{\mathbb{Z}}^{\times}$-invariant, we may view $w$ as function on $G(k^{\mathrm{sep}})$ in a canonical way; namely the value of $w$ via the induced map $G(-1)(k^{\mathrm{sep}})\to G(k^{\mathrm{sep}})$ from Lemma 3.3 is independent of the choice of primitive root of unity. This always holds if $k$ contains no non-trivial roots of unity and is imposed in the literature in many cases, for example [91, §2.1] and [40, §4.2]. It gives simpler formulae for height functions, but it is not relevant to our theory in general.
@@
What we call a weight function, Darda and Yasuda refer to as a raising function. We prefer the former terminology, since crucial to our framework will be the non-trivial conjugacy classes which minimise $w$, which can conveniently be referred to as the minimal weight conjugacy classes.
@@ 3.5. Iitaka fibration
An effective orbifold line bundle $L$ is called rigid if $\kappa(L)=0$, equivalently the conjugacy classes $\mathcal{A}(L)$ generate the entire group $G$. We call $L$ adjoint rigid (or fair) if $\operatorname{ad}(L)$ is rigid; equivalently if the (non-identity) minimal weight conjugacy classes $\mathcal{M}(L):=\mathcal{A}(\operatorname{ad}(L))$ generate $G$ (see Definition 3.4).
@@
Wood defined the notion of fair heights in [91, §2.1] in the case that $G$ is abelian. Fairness (or adjoint rigidity) is actually a property of the underlying orbifold line bundle and not the choice of height; our definition is strictly weaker than Wood’s in the abelian case (see §10.7.4). Our condition in fact appears in Alberts’ paper [7, §7.6], however Alberts does not call such height functions fair since an example is included in the abelian case where the leading constant in Malle’s conjecture need not satisfy the Malle–Bhargava heuristics. However this example is not actually pathological and can be explained by a Brauer–Manin obstruction; again see §10.7.4 for more details. Thus we believe that our definition of fairness is the correct one.
@@ 3.6. Effective cone constant
From now on to define the effective cone constant, we assume that $L$ is fair. We let $\operatorname{Eff}_{G}(L)$ be the image of $\operatorname{Eff}_{G}$ inside the real vector space $\operatorname{Pic}^{\mathrm{orb}}_{L}BG\otimes\mathbb{R}$.
@@
Let $L$ be a fair orbifold line bundle. Then
@@ 3.7. Breaking cocycles
In the case that $L$ is fair, the condition $\langle\mathcal{M}(L)\rangle\subseteq H$ implies that $BH\cong BG$, which is a trivial case. For unfair $L$ however, this condition covers more cases: for example any fibre of the Iitaka fibration will satisfy $b(k,f^{*}L)\geq b(k,L)$. Not all such cocycles should be removed (indeed, they cover $BG(k)$), but only those which give a strict inequality should be removed.
@@
The following summarises the results of this section and is a generalisation of Theorem 1.12. Theorem 3.40(1) should be viewed as an analogue of [57, Thm. 1.4] in the setting of $BG$. More than this however, we are able to give an explicit description of the breaking thin set, something which is currently lacking the setting of Manin’s conjecture.
@@ 4.1. The stack $\operatorname{Spec}\mathcal{O}_{\sqrt[n]{v}}$
Unfolding the definition of $i_{\pi}$ we can describe it as follows. Let $S\to\operatorname{Spec}\mathbb{F}$ be an $\mathbb{F}$-scheme. A point of $B\mu_{n}(S)$ corresponds to a $\mu_{n}$-torsor, which by Kummer theory [84, Tag 03PK] is the same as a pair $(\mathcal{L},\eta)$, where $\mathcal{L}$ is a line bundle on $S$ and $\eta:\mathcal{L}^{n}\cong\mathcal{O}_{S}$ is an isomorphism. The map $i_{\pi}$ sends this to the pair $(\mathcal{L},\pi\cdot\eta)$ where $\pi\cdot\eta:\mathcal{L}^{n}\cong\mathcal{O}_{S}\xrightarrow{s\to\pi\cdot s}\pi\cdot\mathcal{O}_{S}$ is an isomorphism, (that this defines an element of $\operatorname{Spec}{\mathcal{O}_{\sqrt[n]{v}}}$ follows from the definition.)
@@ 4.2.1. Twisting torsors
We first recall some facts on twisting torsors. The following is based upon [83, p. 20–22]. Let $S$ be a scheme, $G$ a group scheme over $S$ acting on the left on an $S$-scheme $X$. Let $a:P\to S$ be a right $G$-torsor. It is also a left-torsor under the inner twist $G_{a}$ [83, p. 20 Ex. 1] of $G$ by $a$.
@@ 4.2.2. Quotient stacks
Recall [84, Tag 04UV] that for any scheme $T\to S$ the groupoid $[X/G](T)$ consists of a pair $(a,f)$ of a right $G$-torsor $a:P\to T$ equipped with a map $f:P\to X$ such that $f(p\cdot g^{-1})=g\cdot f(p)$. Morphisms $(a,f)\to(a^{\prime},f^{\prime})$ are given by a commutative triangle
@@ 4.2.3. Maps of quotient stacks
This defines a map from the groupoid scheme $G\times X\rightrightarrows X$ to $G^{\prime}\times X^{\prime}\rightrightarrows X^{\prime}$ in the sense of [84, Tag 0230] and thus induces a map on quotient stacks $[X/G]\to[X^{\prime}/G^{\prime}]$.
@@ 4.5. The case of $BG$
Let us now recall the structure of $\Gamma_{K}$, cf. [84, Tag 09E3]. Let $\Gamma_{k}^{\text{tame}}$ be Galois group of the maximal tamely ramified extension of $K$. The kernel of $\Gamma_{K}\to\Gamma_{k}^{\text{tame}}$ is a pro-$p$-group. Let $I_{K}^{\text{tame}}\subseteq\Gamma_{k}^{\text{tame}}$ be the tame inertia subgroup. We have a short exact sequence
@@
Moreover, by [84, Tag 09EE] there is a canonical isomorphism
@@ 4.6. Stacky Hensel’s Lemma
Let $\mathcal{X}$ be a smooth proper stack over a henselian DVR $\mathcal{O}$. Then the usual version of Hensel’s lemma says that $\mathcal{X}(\mathcal{O})\to\mathcal{X}(\mathbb{F})$ is surjective (see e.g. [61, Lem. 4.2] for the case where $\mathcal{O}$ is complete). If $\mathcal{X}$ is a scheme then this automatically applies to $\mathcal{X}(K)=\mathcal{X}(\mathcal{O})$. However in general we have $\mathcal{X}(K)\neq\mathcal{X}(\mathcal{O})$ due to the failure of the usual valuative criterion for properness for stacks. The following can be viewed as generalisation of Hensel’s lemma for proper 00-dimensional stacks in light of this failure. It says that $\mathcal{X}(K)$ can still be related to the $\mathbb{F}$-points of a stack, not $\mathcal{X}$ itself, but rather the cyclotomic inertia stack. It shows that the cyclotomic inertia stacks plays the role of a smooth proper compactification of $\mathcal{X}$ (even though $\mathcal{X}$ itself is already smooth and proper!). This result highlights the arithmetic significance of the cyclotomic inertia stack and will be used in the proof of our generalisation of Bhargava’s mass formula (Corollary 8.11). The following is a slightly more general version of Theorem 1.5.
@@ 5.2. Points on stacks
If $\mathcal{X}$ is an algebraic space then it has a dense open subscheme [84, Tag 06NH] so we take $K$ the function field of $\mathcal{X}$.
@@
By Chebotarev’s density theorem for schemes [74, Thm. B.9.], the set of closed points $u\in U$ such that $\pi^{-1}(u)$ is connected has positive Dirichlet density [74, Def. B.6]. Moreover, the set of closed points $u\in U$ whose image $v\in\operatorname{Spec}\mathcal{O}_{k,S}$ is such that $\mathbb{F}(u)=\mathbb{F}_{v}$ has Dirichlet density $1$ by [74, Prop. B.8]. It follows that we may choose $u$ satisfying both properties. The second property implies that $u\in U(\mathbb{F}_{v})$ and the first property is exactly what we needed to conclude the proof. ∎
@@ 5.3. Unramified Brauer group
If $\mathcal{O}$ is a henselian DVR with fraction field $K$ and $\hat{\mathcal{O}}$ is the completion with fraction field $\hat{K}$ then $\operatorname{Br}\mathcal{O}=\operatorname{Br}\hat{\mathcal{O}}$ and $\operatorname{Br}K=\operatorname{Br}\hat{K}$ by [30, Cor. 3.4.3, Prop. 7.1.8]. But by Cohen’s structure’s theorem [84, Tag 0C0S] we have $\hat{\mathcal{O}}\cong\mathbb{F}[[t]]$ where $\mathbb{F}$ is the residue field of $\mathcal{O}$. This implies the second part of the lemma. ∎
@@ 5.6.2. Unramified Brauer–Manin pairing
We similarly define $(\prod_{v}X(k_{v}))^{\operatorname{Br_{un}}}$ to be the left kernel of this pairing and have $X(k)\subseteq(\prod_{v}X(k_{v}))^{\operatorname{Br_{un}}}$. This pairing seems to be more fundamental to the theory and can be used to give obstructions to weak approximation. (Here we say that $\mathcal{X}$ satisfies weak approximation if the image of $\mathcal{X}[k]$ in $\prod_{v}\mathcal{X}[k_{v}]$ is dense; some background on this for stacks can be found in [61, §4].) For example, it allows one to recover Wang’s counter-example to Grunwald’s theorem as a Brauer–Manin obstruction (see §7.5). To verify that this pairing is well-defined we require the following result, which is a minor generalistion of Theorem 1.9. For varieties this is well-known and part of Harari’s formal lemma [30, Thm. 13.4.1]. We generalise this result to the case of the partially unramified Brauer group of $BG$ in Theorem 7.4.
@@ 5.7. Brauer groups of $\mu_{n}$-gerbes
For cup products there is an explicit description of $\operatorname{\partial}_{\mathcal{X}}$; the following argument is inspired by the proof of [83, Thm. 4.1.1].
@@
The map $\operatorname{\partial}_{\mathcal{X},n}$ is a boundary map of the Leray spectral sequence of $f:(B\mu_{n})_{\mathcal{X}}\to\mathcal{X}$. The Leray spectral sequence has cup products [85, Cor. 8.8] which implies that $\operatorname{\partial}_{\mathcal{X},n}(\alpha\cup[\pi])$ is the cup product of $\alpha\in\operatorname{H}^{1}(\mathcal{X},\hat{T})$ and the image of $[\pi]$ under the boundary map $d:\operatorname{H}^{1}((B\mu_{n})_{\mathcal{X}},T)\to\operatorname{H}^{0}(\mathcal{X},\operatorname{R}^{1}T)$. The same computation as in Lemma 5.24 shows that $\operatorname{R}^{1}f_{*}T\cong\operatorname{Hom}(\mu_{n},T)$.
@@ 5.8. Brauer groups of root stacks
This is a minor generalization of [62, Thm. 1]; at first sight there seems to be a sign difference, but this is due to different sign conventions for the Leray spectral sequence.
@@ 5.10. A geometric interpretation of the residue
An element of $[\pi]\in\operatorname{H}^{2}(BG,T)$ corresponds to a $T$-gerbe $\pi:\mathcal{T}\to BG$ as in [69, §12.2.7]. The condition that $[\pi]\in\operatorname{H}_{1}^{2}(BG,T)$ means that the gerbe is locally neutral on $\mathcal{X}$ (rather than just locally neutral on $BG$), i.e. locally on $\mathcal{X}$ of the form $BT\times G$. In particular, $\mathcal{T}$ is a gerbe with banded by an abelian group scheme $A$ over $\mathcal{X}$. The map $\pi:\mathcal{T}\to BG$ induces a map $A\to G$ on bands. Moreover, this map is surjective and has kernel $T$, both of these facts are checked locally on $\mathcal{X}$. In particular, $A$ defines an element of $\operatorname{Ext}^{1}_{\mathcal{X}}(G,T)$. Moreover, this extensions splits locally on $\mathcal{X}$ and by the Leray spectral sequence $\operatorname{H}^{p}(\mathcal{X},\underline{\operatorname{Ext}}^{q}(G,T))\implies\operatorname{Ext}^{p+q}(G,T)$ thus lives as a class $[A]\in\operatorname{H}^{1}(\mathcal{X},\underline{\operatorname{Hom}}(G,T))$.
@@
Applying the functor $\tau_{\leq 0}Rf_{*}$ and using the fact that $\tau_{\leq 0}Rf_{*}\tau_{\leq 0}R\pi_{*}T[1]\cong\tau_{\leq 0}R(f\circ\pi)T[1]$ by [84, Tag 015M] we get a triangle
@@
The description of spectral sequences in terms of derived functors (as explained in e.g. [83, pp. 67–68]) shows that $r([\pi])$ is equal to the image of $\text{id}_{T}$ under the composition of the maps $\underline{\operatorname{Hom}}(T,T)\to\tau_{\leq 1}Rf_{*}T[2]\to R^{1}f_{*}T[1]\cong\underline{\operatorname{Hom}(G,T)}[1]$.

### v3

@@ 3.2. Orbifold Picard group of $BG$
Condition (2) says that $\zeta^{w(\gamma)}=\chi(\gamma(\zeta))$ for all $\gamma\in G(-1)(k^{\mathrm{sep}})$ and all $\zeta\in\mu_{n}$ with $n$ not divisible by the characteristic of $k$; via a choice of isomorphism from Lemma 3.3 this becomes $\zeta^{w(g)}=\chi(g)$ for all $g\in G(k^{\mathrm{sep}})$ and all such $n$. In particular $w$ takes integer values if and only if $\chi$ is trivial.
@@
If $w$ is $\widehat{\mathbb{Z}}^{\times}$-invariant, we may view $w$ as function on $G(k^{\mathrm{sep}})$ in a canonical way; namely the value of $w$ via the induced map $G(-1)(k^{\mathrm{sep}})\to G(k^{\mathrm{sep}})$ from Lemma 3.3 is independent of the choice of primitive root of unity. This always holds if $k$ contains no non-trivial roots of unity and is imposed in the literature in many cases, for example [93, §2.1] and [40, §4.2]. It gives simpler formulae for height functions, but it is not relevant to our theory in general.
@@
What we call a weight function, Darda and Yasuda refer to as a raising function. We use the former terminology, since crucial to our framework will be the non-trivial conjugacy classes which minimise $w$, which can conveniently be referred to as the minimal weight conjugacy classes.
@@ 3.5. Iitaka fibration
An effective orbifold line bundle $L$ is called rigid if $\kappa(L)=0$, equivalently the conjugacy classes $\mathcal{A}(L)$ generate the entire group $G$. We call $L$ adjoint rigid (or balanced) if $\operatorname{ad}(L)$ is rigid; equivalently if the (non-identity) minimal weight conjugacy classes $\mathcal{M}(L):=\mathcal{A}(\operatorname{ad}(L))$ generate $G$ (see Definition 3.4).
The terminology balanced is used to conjure the image that there are enough minimal weight elements to balance a plate on top of; as opposed to the concentrated case from [6] where the minimal weight elements are too clumped together to balance a plate on. We use the terminology balanced line bundle in a different way to the paper [57]; however this terminology has subsequently been replaced in the literature by the “breaking” terminology from [58].
@@
Wood defined the notion of fair heights in [93, §2.1] in the case that $G$ is abelian. Our definition of balanced (or adjoint rigidity) is strictly weaker than Wood’s in the abelian case (see §10.7.4), and is a property of the underlying orbifold line bundle and not the choice of height.
Our condition also appears in Alberts’ paper [7, §7.6], however Alberts does not call such height functions fair since an example is included in the abelian case where the leading constant in Malle’s conjecture need not satisfy the Malle–Bhargava heuristics. However this example is not as pathological as one might first think, as it can be explained by a Brauer–Manin obstruction; again see §10.7.4 for more details. Thus we believe that our definition of balancedness is the correct one to encapsulate a measure-theoretic interpretation of the leading constant in Malle’s conjecture.
@@ 3.6. Effective cone constant
From now on to define the effective cone constant, we assume that $L$ is balanced. We let $\operatorname{Eff}_{G}(L)$ be the image of $\operatorname{Eff}_{G}$ inside the real vector space $\operatorname{Pic}^{\mathrm{orb}}_{L}BG\otimes\mathbb{R}$.
@@
Let $L$ be a balanced orbifold line bundle. Then
@@ 3.7. Breaking cocycles
In the case that $L$ is balanced, the condition $\langle\mathcal{M}(L)\rangle\subseteq H$ implies that $BH\cong BG$, which is a trivial case. For unbalanced $L$ however, this condition covers more cases: for example any fibre of the Iitaka fibration will satisfy $b(k,f^{*}L)\geq b(k,L)$. Not all such cocycles should be removed (indeed, they cover $BG(k)$), but only those which give a strict inequality should be removed.
@@
The following summarises the results of this section and is a generalisation of Theorem 1.12. Theorem 3.40(1) should be viewed as an analogue of [58, Thm. 1.4] in the setting of $BG$. More than this however, we are able to give an explicit description of the breaking thin set, something which is currently lacking the setting of Manin’s conjecture.
@@ 4.1. The stack $\operatorname{Spec}\mathcal{O}_{\sqrt[n]{v}}$
Unfolding the definition of $i_{\pi}$ we can describe it as follows. Let $S\to\operatorname{Spec}\mathbb{F}$ be an $\mathbb{F}$-scheme. A point of $B\mu_{n}(S)$ corresponds to a $\mu_{n}$-torsor, which by Kummer theory [86, Tag 03PK] is the same as a pair $(\mathcal{L},\eta)$, where $\mathcal{L}$ is a line bundle on $S$ and $\eta:\mathcal{L}^{n}\cong\mathcal{O}_{S}$ is an isomorphism. The map $i_{\pi}$ sends this to the pair $(\mathcal{L},\pi\cdot\eta)$ where $\pi\cdot\eta:\mathcal{L}^{n}\cong\mathcal{O}_{S}\xrightarrow{s\to\pi\cdot s}\pi\cdot\mathcal{O}_{S}$ is an isomorphism, (that this defines an element of $\operatorname{Spec}{\mathcal{O}_{\sqrt[n]{v}}}$ follows from the definition.)
@@ 4.2.1. Twisting torsors
We first recall some facts on twisting torsors. The following is based upon [85, p. 20–22]. Let $S$ be a scheme, $G$ a group scheme over $S$ acting on the left on an $S$-scheme $X$. Let $a:P\to S$ be a right $G$-torsor. It is also a left-torsor under the inner twist $G_{a}$ [85, p. 20 Ex. 1] of $G$ by $a$.
@@ 4.2.2. Quotient stacks
Recall [86, Tag 04UV] that for any scheme $T\to S$ the groupoid $[X/G](T)$ consists of a pair $(a,f)$ of a right $G$-torsor $a:P\to T$ equipped with a map $f:P\to X$ such that $f(p\cdot g^{-1})=g\cdot f(p)$. Morphisms $(a,f)\to(a^{\prime},f^{\prime})$ are given by a commutative triangle
@@ 4.2.3. Maps of quotient stacks
This defines a map from the groupoid scheme $G\times X\rightrightarrows X$ to $G^{\prime}\times X^{\prime}\rightrightarrows X^{\prime}$ in the sense of [86, Tag 0230] and thus induces a map on quotient stacks $[X/G]\to[X^{\prime}/G^{\prime}]$.
@@ 4.5. The case of $BG$
Let us now recall the structure of $\Gamma_{K}$, cf. [86, Tag 09E3]. Let $\Gamma_{k}^{\text{tame}}$ be Galois group of the maximal tamely ramified extension of $K$. The kernel of $\Gamma_{K}\to\Gamma_{k}^{\text{tame}}$ is a pro-$p$-group. Let $I_{K}^{\text{tame}}\subseteq\Gamma_{k}^{\text{tame}}$ be the tame inertia subgroup. We have a short exact sequence
@@
Moreover, by [86, Tag 09EE] there is a canonical isomorphism
@@ 4.6. Stacky Hensel’s Lemma
Let $\mathcal{X}$ be a smooth proper stack over a henselian DVR $\mathcal{O}$. Then the usual version of Hensel’s lemma says that $\mathcal{X}(\mathcal{O})\to\mathcal{X}(\mathbb{F})$ is surjective (see e.g. [62, Lem. 4.2] for the case where $\mathcal{O}$ is complete). If $\mathcal{X}$ is a scheme then this automatically applies to $\mathcal{X}(K)=\mathcal{X}(\mathcal{O})$. However in general we have $\mathcal{X}(K)\neq\mathcal{X}(\mathcal{O})$ due to the failure of the usual valuative criterion for properness for stacks. The following can be viewed as generalisation of Hensel’s lemma for proper 00-dimensional stacks in light of this failure. It says that $\mathcal{X}(K)$ can still be related to the $\mathbb{F}$-points of a stack, not $\mathcal{X}$ itself, but rather the cyclotomic inertia stack. It shows that the cyclotomic inertia stacks plays the role of a smooth proper compactification of $\mathcal{X}$ (even though $\mathcal{X}$ itself is already smooth and proper!). This result highlights the arithmetic significance of the cyclotomic inertia stack and will be used in the proof of our generalisation of Bhargava’s mass formula (Corollary 8.11). The following is a slightly more general version of Theorem 1.5.
@@ 5.2. Points on stacks
If $\mathcal{X}$ is an algebraic space then it has a dense open subscheme [86, Tag 06NH] so we take $K$ the function field of $\mathcal{X}$.
@@
By Chebotarev’s density theorem for schemes [76, Thm. B.9.], the set of closed points $u\in U$ such that $\pi^{-1}(u)$ is connected has positive Dirichlet density [76, Def. B.6]. Moreover, the set of closed points $u\in U$ whose image $v\in\operatorname{Spec}\mathcal{O}_{k,S}$ is such that $\mathbb{F}(u)=\mathbb{F}_{v}$ has Dirichlet density $1$ by [76, Prop. B.8]. It follows that we may choose $u$ satisfying both properties. The second property implies that $u\in U(\mathbb{F}_{v})$ and the first property is exactly what we needed to conclude the proof. ∎
@@ 5.3. Unramified Brauer group
If $\mathcal{O}$ is a henselian DVR with fraction field $K$ and $\hat{\mathcal{O}}$ is the completion with fraction field $\hat{K}$ then $\operatorname{Br}\mathcal{O}=\operatorname{Br}\hat{\mathcal{O}}$ and $\operatorname{Br}K=\operatorname{Br}\hat{K}$ by [30, Cor. 3.4.3, Prop. 7.1.8]. But by Cohen’s structure’s theorem [86, Tag 0C0S] we have $\hat{\mathcal{O}}\cong\mathbb{F}[[t]]$ where $\mathbb{F}$ is the residue field of $\mathcal{O}$. This implies the second part of the lemma. ∎
@@ 5.6.2. Unramified Brauer–Manin pairing
We similarly define $(\prod_{v}X(k_{v}))^{\operatorname{Br_{un}}}$ to be the left kernel of this pairing and have $X(k)\subseteq(\prod_{v}X(k_{v}))^{\operatorname{Br_{un}}}$. This pairing seems to be more fundamental to the theory and can be used to give obstructions to weak approximation. (Here we say that $\mathcal{X}$ satisfies weak approximation if the image of $\mathcal{X}[k]$ in $\prod_{v}\mathcal{X}[k_{v}]$ is dense; some background on this for stacks can be found in [62, §4].) For example, it allows one to recover Wang’s counter-example to Grunwald’s theorem as a Brauer–Manin obstruction (see §7.5). To verify that this pairing is well-defined we require the following result, which is a minor generalistion of Theorem 1.9. For varieties this is well-known and part of Harari’s formal lemma [30, Thm. 13.4.1]. We generalise this result to the case of the partially unramified Brauer group of $BG$ in Theorem 7.4.
@@ 5.7. Brauer groups of $\mu_{n}$-gerbes
For cup products there is an explicit description of $\operatorname{\partial}_{\mathcal{X}}$; the following argument is inspired by the proof of [85, Thm. 4.1.1].
@@
The map $\operatorname{\partial}_{\mathcal{X},n}$ is a boundary map of the Leray spectral sequence of $f:(B\mu_{n})_{\mathcal{X}}\to\mathcal{X}$. The Leray spectral sequence has cup products [87, Cor. 8.8] which implies that $\operatorname{\partial}_{\mathcal{X},n}(\alpha\cup[\pi])$ is the cup product of $\alpha\in\operatorname{H}^{1}(\mathcal{X},\hat{T})$ and the image of $[\pi]$ under the boundary map $d:\operatorname{H}^{1}((B\mu_{n})_{\mathcal{X}},T)\to\operatorname{H}^{0}(\mathcal{X},\operatorname{R}^{1}T)$. The same computation as in Lemma 5.24 shows that $\operatorname{R}^{1}f_{*}T\cong\operatorname{Hom}(\mu_{n},T)$.
@@ 5.8. Brauer groups of root stacks
This is a minor generalization of [63, Thm. 1]; at first sight there seems to be a sign difference, but this is due to different sign conventions for the Leray spectral sequence.
@@ 5.10. A geometric interpretation of the residue
An element of $[\pi]\in\operatorname{H}^{2}(BG,T)$ corresponds to a $T$-gerbe $\pi:\mathcal{T}\to BG$ as in [71, §12.2.7]. The condition that $[\pi]\in\operatorname{H}_{1}^{2}(BG,T)$ means that the gerbe is locally neutral on $\mathcal{X}$ (rather than just locally neutral on $BG$), i.e. locally on $\mathcal{X}$ of the form $BT\times G$. In particular, $\mathcal{T}$ is a gerbe with banded by an abelian group scheme $A$ over $\mathcal{X}$. The map $\pi:\mathcal{T}\to BG$ induces a map $A\to G$ on bands. Moreover, this map is surjective and has kernel $T$, both of these facts are checked locally on $\mathcal{X}$. In particular, $A$ defines an element of $\operatorname{Ext}^{1}_{\mathcal{X}}(G,T)$. Moreover, this extensions splits locally on $\mathcal{X}$ and by the Leray spectral sequence $\operatorname{H}^{p}(\mathcal{X},\underline{\operatorname{Ext}}^{q}(G,T))\implies\operatorname{Ext}^{p+q}(G,T)$ thus lives as a class $[A]\in\operatorname{H}^{1}(\mathcal{X},\underline{\operatorname{Hom}}(G,T))$.
@@
Applying the functor $\tau_{\leq 0}Rf_{*}$ and using the fact that $\tau_{\leq 0}Rf_{*}\tau_{\leq 0}R\pi_{*}T[1]\cong\tau_{\leq 0}R(f\circ\pi)T[1]$ by [86, Tag 015M] we get a triangle
@@
The description of spectral sequences in terms of derived functors (as explained in e.g. [85, pp. 67–68]) shows that $r([\pi])$ is equal to the image of $\text{id}_{T}$ under the composition of the maps $\underline{\operatorname{Hom}}(T,T)\to\tau_{\leq 1}Rf_{*}T[2]\to R^{1}f_{*}T[1]\cong\underline{\operatorname{Hom}(G,T)}[1]$.